OpenStudy (anonymous):

Please help me!!! Two opposite sides of a square with length x are increased by 3 units each and the other two opposite sides are decreased by 3 units each. If the area of the resulting rectangle is 27 unit2, what would be the width of the rectangle?

OpenStudy (tkhunny):

1) It depends on which pair of sides you are calling the width. 2) Area of square with sides of length x, Area = x*x 3) Area of defined rectangle, Area = (x+3)*(x-3) = 27 Where does that leave us?

OpenStudy (anonymous):

You solve for x and you will get the answer.

OpenStudy (blank ):

(x+3)*(x-3) = 27 (6+3)*(6-3) = 27 (9)*(3) = 27 Looks like the longest side of the rectangle would be 9 units.

OpenStudy (tkhunny):

Wow! Of course! Just pull it out of your hat. Wave your magic wand and it happens! (x+3)*(x-3) = 27 x^2 - 9 = 27 x^2 - 36 = 0 (x+6)(x-6) = 0 x = -6 or x = 6 As this is a real rectangle, x = -6 is no good. Throw it out! x = 6